Hello I'm having trouble solving a equation containing two radicals. I keep getting the wrong answer even though I isolate one radical, square it, isolate the second, and square that.
The equation is
sqrt(5x-7) + sqrt(2x+8) +2 = 0
Thank you for the help in advance!

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

sqrt(5x-7) + sqrt(2x+8) +2 = 0
-----------------------------
If you post your work, we can check it.

Answer by stanbon(75887)

sqrt(5x-7) + sqrt(2x+8) +2 = 0
-------
sqrt(5x-7) = -[sqrt(2x+8)+2)
----
Square both sides:
5x-7 = 2x+8 + 4sqrt(2x+8)+4
----
3x-19 = 4sqrt(2x+8)
----
Square both sides:
9x^2 - 114x + 361 = 16(2x+8)
9x^2 - 114x + 361 = 32x + 128
----
9x^2 -146x + 233 = 0
----
Use the Quadratic formula to get:
x = 14.428
================================
cheers,
Stan H.